OpenStudy (anonymous):

please help solve using the square root property. 2x^2-4=-166

OpenStudy (anonymous):

2x^2=-162 x^2=-81 x=-9

OpenStudy (anonymous):

x=+and-9

OpenStudy (radar):

Always check your answer for validity. In this case: 2(-9)^2 -4 = -166 2(81) - 4 = -166 162 - 4 = -166 158 = -166 Ooops It is not a -9, well lets check further is it the +9 2(9)^2 - 4 = -166 162 - 4 = -166 again 158 = -166 OOoops again. Neither is correct!! What went wrong??? Lets work this problem again 2x^2 - 4 = -166 Lets add 4 to both sides of the equal sign (a legal operation) 2x^2 = =-166 + 4 = -162 2x^2 = -162 Now lets divide both sides by 2.........another legal operation. x^2 = -81 What now? Looks like all we need to do is take the square root of both sides. Not so fast...note the negative value on the right the -81. Pay attention. There is no real number that when squared that will equal -81 !! What is one to do to solve this ??? Use the imaginary number "i" That is the number when squared equals a -1, or i^2=-1, better yet i = the square root of -1. So the problem is now like this:

OpenStudy (radar):

sqrt x^2 = sqrt -81 sqrt x^2 = sqrt [81 (-1)] sqrt x^2 = sqrt [81 (i^2)] x = 9i Now lets validate (if possible) this answer 2x^2 -4 = -166 (the given problem) 2(9i)^2 - 4 = -166 (2)(81)(i^2) -4 = - 166 Now remember i^2 is a -1 (2)(81)(-1) - 4 = -166 162(-1) -4 = -166 -162 -4 = -166 -166 = -166 Leaving it up to you to prove that -9i is also an answer.

OpenStudy (radar):

@jmedina163 @donmath do you understand?